OpenStudy (anonymous):

Find the domain: [1 / sqroot of 2x+3] + (sqroot of 6-4x)

OpenStudy (zzr0ck3r):

so we have a fraction with a square root on the bottom so what ever is in the square root must be greater than 0 and for the square root on the right, we need the stuff on the inside to be >=0 so we need the intersection of 2x+3>0 and 6-4x>=0 x>-3/2 and x<=3/2 so we need -3/2<x<=3/2

OpenStudy (zzr0ck3r):

does this make sense?

OpenStudy (anonymous):

Yes! Thank you so much!!!

OpenStudy (zzr0ck3r):

np